OpenStudy (anonymous):

log base 4* (1/16) = -2?

OpenStudy (psymon):

\[\log_{b} M = x\] Same as: \[b ^{x} = M\] So in this case that would be correct because 4^(-2) does = 1/16

OpenStudy (anonymous):

can I ask you to help with another?

OpenStudy (psymon):

Go for it.

OpenStudy (anonymous):

Simplify... (Sin^3(x)+Cos^2(x)Sin(x))/(Sin^2(x)) = Sin(x)(1+Cos^2(x))

OpenStudy (anonymous):

is that correct?

OpenStudy (psymon):

Well, let's see.

OpenStudy (psymon):

I would say that it is not. Try factoring out a sinx from the numerator and see where that takes you.

OpenStudy (psymon):

\[\frac{ \sin ^{3}(x) + \cos ^{2}(x)\sin(x) }{ sinx }\] This can become: \[\frac{ sinx(\sin ^{2}x + \cos ^{2}x) }{ sinx }\]

OpenStudy (anonymous):

is it 1?

OpenStudy (psymon):

Yep ^_^

OpenStudy (anonymous):

you are the best :)